2.4/4= x/7
find the
unknown number round to nearest hundredth when needed

Answer by checkley75(3666)

2.4/4=X/7 NOW CROSS MULTIPLY
4X=2.4*7
4X=16.8
X=16.8/4
X=4.2 ANSWER.